# Lines of Thought in Large Language Models

### v2

@@ Abstract
Large Language Models achieve next-token prediction by transporting a vectorized piece of text (prompt) across an accompanying embedding space under the action of successive transformer layers. The resulting high-dimensional trajectories realize different contextualization, or ‘thinking’, steps, and fully determine the output probability distribution. We aim to characterize the statistical properties of ensembles of these ‘lines of thought.’ We observe that independent trajectories cluster along a low-dimensional, non-Euclidean manifold, and that their path can be well approximated by a stochastic equation with few parameters extracted from data. We find it remarkable that the vast complexity of such large models can be reduced to a much simpler form, and we reflect on implications.
@@ 1 Introduction
How does a large language model (LLM) think? In other words, how does it abstract the prompt “Once upon a time, a facetious” to suggest adding, e.g., “transformer”, and, by repeating the operation, continue on to generate a respectable fairy tale à la Perrault? What we know is by design. A piece of text is mapped into a set of high-dimensional vectors, which are then transported across their embedding (latent) space through successive transformer layers (Vaswani et al., 2017), each allegedly distilling different syntactic, semantic, informational, contextual aspects of the input (Valeriani et al., 2023; Song & Zhong, 2024). The final position is then projected onto an embedded vocabulary to create a probability distribution about what the next word should be. Why these vectors land where they do eludes human comprehension due to the concomitant astronomical numbers of arithmetic operations which, taken individually, do nothing, but collectively confer the emergent ability of language.
@@
    We introduce a stochastic model to describe trajectory ensembles with few parameters.
@@ 2 Methods
This section describes our algorithm for generating and analyzing an ensemble of tokens trajectories in the latent space of LLMs.
@@
We form trajectories by collecting the successive vector outputs, within the latent space, after each transformer layer (hidden_states). For conciseness, we identify layer number with a notional ‘time’, $t$. Even though all embedded tokens of a prompt voyage across the latent space, only the vector corresponding to the last token is eventually projected onto $\mathcal{V}$ to form the logits. Hence, here, we only consider the trajectory of this last (or ‘pilot’) token. The trajectory ${\bm{M}}_{k}$ of sentence $k$’s pilot is the sequence of $24$ successive time positions $\{{\bm{x}}_{k}(1),{\bm{x}}_{k}(2),\ldots,{\bm{x}}_{k}(24)\}$, concatenated as a column matrix (Algorithm 1).
@@
6:     ${\bm{E}}(0)={\bm{E}}_{S}+{\bm{E}}_{P}$add positional embeddings ${\bm{P}}$}
7:     for $t=1\to 24$ do
8:        ${\bm{E}}(t+1)=\text{TransformerLayer}_{t}({\bm{E}}(t))$update embeddings through transformer layer}
9:        ${\bm{x}}(t+1)={\bm{E}}(t+1)[:,\text{end}]$extract last token representation}
10:        ${\bm{M}}[:,t+1]={\bm{x}}(t+1)$save trajectory array}
@@
13:  Output: Final embeddings $x(t+1)$ for all pseudo-sentences
@@
The latent space is spanned by a Cartesian basis, i.e., the orthogonal set of one-hot (unit) vectors =$\mathcal{E}=\{{\bm{e}}_{i}\}_{i=1\dots D}$. [^7] With a 1 in $i^{\mathrm{th}}$ position, 0 elsewhere. Additionally, we will often refer to the bases $\mathcal{U}(t)=\{{\bm{u}}_{i}^{(t)}\}_{i=1\dots D}$ formed by the left-singular vectors of the singular value decomposition (SVD) of the $D\times N_{\mathrm{s}}$ matrix after layer $t$: ${\bm{M}}={\bm{U}}\mathbf{\Sigma}{\bm{V}}^{\top}$, with ${\bm{M}}_{:,k}(t)={\bm{x}}_{k}(t)$. Vectors ${\bm{u}}_{i}$ are organized according to their corresponding singular values, $\sigma_{i}$, in descending order. Note that because trajectory clusters evolve over time there are 24 distinct bases.
@@ 3.1 Lines of thought cluster along similar pathways
We first observe in Fig. 1a that pilot trajectories tend to cluster together, instead of producing an isotropic and homogeneous filling of $\mathcal{S}$. Indeed, LoTs for different, independent pseudo-sentences follow a common path (forming bundles), despite individual variability. Specifically, there exist directions with significant displacement relative to the spread (mean over standard deviation). In addition, positions at different times form distinct clusters, as shown in Fig. 6.
Properly visualizing these trajectories is difficult, due to the high dimensionality of $\mathcal{S}$. Because the Cartesian axes, ${\bm{e}}_{i}$, are unlikely to align with trajectories meaningful directions, we seek relevant alternative bases, informed by the data. After each layer $t$, we perform the singular value decomposition of the matrix formed by concatenating the ${\bm{x}}_{k}(t)$ to obtain a basis $\mathcal{U}(t)$ aligned with the data’s intrinsic directions. In the following, we leverage these time-dependent singular vectors and values to investigate ensemble dynamics and low-dimensional structures.
@@ 3.2 Lines of thought follow a low-dimensional manifold
The fast decay of the singular value magnitudes seen in Fig. 2b suggests that LoTs may be described by a lower-dimensional subspace. But how many dimensions are relevant? Singular values relate to ensemble variance along their corresponding directions. Since the embedding space is high-dimensional, however, the curse of dimensionality looms, hence the significance of Euclidean distances crumbles. To circumvent this limitation, we consider a more practical metric: how close to the original output distribution on the vocabulary does a reduction in dimensionality get us?
@@
where the ${\bm{u}}_{i}^{(t)}$’s are organized by descending order of their corresponding singular values. By default $K=D$, and the true output distribution ${\bm{p}}^{\mathcal{V}}$ is obtained. Now, we examine what happens when, instead of passing the full basis set, we truncate it, after each layer, to keep only the first $K<D$ principal components. We compare the resulting output distribution, $\mathbf{p}^{\mathcal{V}}_{K}$ to the true distribution $\mathbf{p}^{\mathcal{V}}$ using KL divergence $D_{\mathrm{KL}}(\mathbf{p}^{\mathcal{V}}_{K}\|\mathbf{p}^{\mathcal{V}})$. In Fig. 2c, we see that most of the true distribution is recovered when keeping only about $K_{0}=256$, or $25\%$, of the principal components. In other words, for the purpose of next-token prediction, LoTs are quasi-256-dimensional.
@@
Figure 2: (a) Angle between the first 4 singular vectors at $(t_{1},t_{2})$, $\arccos({\bm{u}}_{i}^{(t_{1})}\cdot{\bm{u}}_{i}^{(t_{2})})$, for $i=\{1,2,3,4\}$ (top-left, top-right, bottom-left, bottom-right, respectively). (b) Singular values for $t=1,\dots,24$ (blue to red). Clusters stretch more and more after each layer. The leading singular values, $\sigma_{1}(t)$, have been omitted for clarity. (c) Average (over all trajectories) KL divergence between reduced dimensionality trajectories output and true output distributions, as the dimensionality $K$ is increased. The red dashes line shows the average KL divergence between unrelated distributions (baseline for dissimilar distributions).
@@ 3.3 Linear approximation of trajectories
Examination of the singular vectors and values at each time step indicates that LoT bundles rotate and stretch smoothly after passing through each layer (Fig. 2). This suggests that token trajectories could be approximated by the linear transformations described by the ensemble, and extrapolated accordingly, from an initial time $t$ to a later time $t+\tau$. Evidently, it is improbable that a transformer layer could be replaced by a mere linear transformation. We rather hypothesize that, in addition to this deterministic average path, a token’s location after layer $t+\tau$ will depart from its linear approximation from $t$ by an unknown component ${\bm{w}}(t,\tau)$.888 We emphasize that trajectories are completely deterministic; the uncertainty (or stochasticity) introduced here accounts only for the loss of information of considering the token without its prompt. We propose the following model:
@@
where ${\bm{x}}(t)$ is the pilot token’s position in the Cartesian basis, and ${\bm{R}},\mathbf{\Lambda}$ are rotation (orthonormal) and stretch (diagonal) matrices, respectively. Eq. 1 formalizes the idea that, to approximate ${\bm{x}}(t+\tau)$, given ${\bm{x}}(t)$, we first project ${\bm{x}}$ in the ensemble intrinsic basis at $t$ (${\bm{R}}^{\top}{\bm{x}}$), then stretch the coordinates by the amount given by $\mathbf{\Lambda}$, and finally rotate according to how much the singular directions have rotated between $t$ and $t+\tau$, ${\bm{R}}(t+\tau)$ (see also Fig. 7 in Appendix B). Consequently, we can express these matrices as a function of the set of singular vectors (${\bm{U}}$) and values ($\mathbf{\Sigma}$):
@@
Fig. 3 shows the close agreement, at the ensemble level, between the true and extrapolated positions. This is merely a linear approximation as it is similar to assuming that LoT clusters deform like an elastic solid, where each point maintains the same vicinity. The actual coordinates ought to include an additional random component ${\mathbf{w}}(t,\tau)$, which a priori depends on both $t$ and $\tau$.
@@
between true positions ${\bm{x}}$ and linear approximations $\tilde{{\bm{x}}}(t,\tau)={\bm{U}}(t+\tau)\mathbf{\Lambda}(t,\tau){\bm{U}}(t)^{\top}{\bm{x}}(t)$. We investigate the distributions and correlations of $\delta{\bm{x}}(t,\tau)$ across layer combinations $(t,t+\tau)$.
From the data, Fig. 4 shows that, for all $(t,t+\tau)\in\{1,\dots,23\}\times\{t+1,\dots,24\}$, the ensemble of $\delta{\bm{x}}(t,\tau)$ has the following characteristics: 1) it is Gaussian, 2) with zero mean, 3) and variance scaling as $\exp(t+\tau)$. In addition, Fig. 8 shows that the distribution is isotropic, with no evidence of spatial cross-correlations . Hence, we propose:
@@ 3.4 Langevin dynamics for continuous time trajectories
In return, this allows us to interpolate trajectories between transformer layers.101010 These interpolated positions do not hold any interpretive value, but may be insightful for mathematical purposes. Thus, we extend Eq. 1 to a continuous time variable $t$, and write in infinitesimal terms:
@@ 4.2.1 Gibberish
We generate non-language (‘gibberish’) pseudo-sentences by assembling $N$-token sequences of random tokens in the vocabulary, and pass them as input to GPT-2. The resulting trajectories also cluster around a path similar to that of language. However, the two ensembles, language and gibberish, are linearly separable at all layers (see Fig. 10 in Section B.5), indicating that they travel on two distinct, yet adjacent, manifolds.
@@ 4.2.2 Untrained & ablated models
First, we collect trajectories of the Walden ensemble passing through a reinitialized version of GPT-2 (the weights have been reset to a random seed). We observe that while LoTs get transported away from their starting point, the trajectories follow straight, quasi-parallel paths, maintaining their vicinity (see Fig. 10). Furthermore, the model of Eqs. 1 and 2 does not hold; Fig. 11 shows that the variance of $\delta{\bm{x}}$ does not follow the $\exp(t+\tau)$ scaling, and the distributions are far from Gaussian.
Next, we consider an ablated model, where only layers $13$ to $24$ have been reinitialized. When reaching the untrained layers, the trajectories stop and merely diffuse about their $t=12$ location (Fig. 10).
@@ 4.3 Results with other models
We first investigate the Llama 2 7B model (Touvron et al., 2023).121212 Decoder-only, 32 layers, 4096 dimensions; released July 2023 by Meta AI. Remarkably, the pattern of GPT-2 repeats. Token positions at $t+\tau$ can be extrapolated from $t$ by rotation and stretch using the singular vectors and values of the ensemble. The residuals are distributed as those of GPT-2, with ${\textnormal{w}}_{i}(t,\tau)\sim\mathcal{N}(0,\alpha e^{\lambda(t+\tau)})$, see Fig. 12. The values for the parameters $\alpha$ and $\lambda$, however, differ from those of GPT-2 (here, $\alpha\simeq-5.4,\lambda\simeq 0.27$).
@@
Trajectories across the Mistral 7B (v0.1) model (Jiang et al., 2023)131313 Decoder-only, 32 layers, 4096 dimensions; released September 2023 by Mistral AI. also follow the same pattern (Fig. 13). We note, however, that Eq. 4 only holds up until layer 31. It seems as though the last layer is misaligned with the rest of the trajectories, as linear extrapolation produces an error that is much larger than expected.
@@
The last layer anomaly is also apparent for Llama 3.2 1B141414 Decoder-only, 16 layers, 2048 dimensions; released September 2024 by Meta AI. , both in the mean and variance of $\delta{\bm{x}}(t,16)$ (see Fig. 14). However, the rest of the trajectories follows Eq. 1. The same pattern is observed for Llama 3.2 3B151515 Decoder-only, 28 layers, 3072 dimensions; released September 2024 by Meta AI. in Fig. 15.

### v3

@@ Abstract
Large Language Models achieve next-token prediction by transporting a vectorized piece of text (prompt) across an accompanying embedding space under the action of successive transformer layers. The resulting high-dimensional trajectories realize different contextualization, or ‘thinking’, steps, and fully determine the output probability distribution. We aim to characterize the statistical properties of ensembles of these ‘lines of thought.’ We observe that independent trajectories cluster along a low-dimensional, non-Euclidean manifold, and that their path can be well approximated by a stochastic equation with few parameters extracted from data. We find it remarkable that the vast complexity of such large models can be reduced to a much simpler form, and we reflect on implications. Code for trajectory generation, visualization, and analysis is available on Github at https://github.com/rapsar/lines-of-thought.
@@ 1 Introduction
How does a large language model (LLM) think? In other words, how does it abstract the prompt “Once upon a time, a facetious” to suggest adding, e.g., “chatbot”, and, by repeating the operation, continue on to generate a respectable fairy tale à la Perrault? What we know is by design. A piece of text is mapped into a set of high-dimensional vectors, which are then transported across their embedding (latent) space through successive transformer layers (Vaswani et al., 2017), each allegedly distilling different syntactic, semantic, informational, contextual aspects of the input (Valeriani et al., 2023; Song & Zhong, 2024). The final position is then projected onto an embedded vocabulary to create a probability distribution about what the next word should be. Why these vectors land where they do eludes human comprehension due to the concomitant astronomical numbers of arithmetic operations which, taken individually, do nothing, but collectively confer the emergent ability of language.
@@
    We introduce a stochastic model to describe trajectory ensembles with few parameters and extend them to continuous paths.
@@ 2 Methods
This section describes our algorithm for generating and analyzing an ensemble of tokens trajectories in the latent space of LLMs. Our code is provided in the corresponding Github repository (Sarfati et al., ).
@@
We form trajectories by collecting the successive vector outputs, within the latent space, after each transformer layer (hidden_states). For conciseness, we identify layer number with a notional ‘time’, $t$. Even though all embedded tokens of a prompt voyage across the latent space, only the embedding corresponding to the last token form the logits (by projection onto $\mathcal{V}$) for next-token inference. Hence, here, we only consider the trajectory of this last (or ‘pilot’) token. The trajectory ${\bm{M}}_{k}$ of sentence $k$’s pilot is the sequence of $24$ successive time positions $\{{\bm{x}}_{k}(1),{\bm{x}}_{k}(2),\ldots,{\bm{x}}_{k}(24)\}$, concatenated as a column matrix (Algorithm 1).
@@
6:     ${\bm{E}}^{0}={\bm{E}}_{S}+{\bm{E}}_{P}$add positional embeddings ${\bm{P}}$}
7:     for $t=1\to 23$ do
8:        ${\bm{E}}^{t+1}=\text{TransformerLayer}_{t}({\bm{E}}^{t})$update embeddings through transformer layer}
9:        ${\bm{x}}(t+1)={\bm{E}}^{(t+1)}_{:,\text{end}}$extract last token representation}
10:        ${\bm{M}}_{:,t+1}={\bm{x}}(t+1)$save trajectory array}
@@
13:  Output: Final embeddings ${\bm{x}}(t+1)$ for all pseudo-sentences
@@
The latent space is spanned by the Cartesian basis =$\mathcal{E}=\{{\bm{e}}_{i}\}_{i=1\dots D}$ (the orthogonal set of one-hot unit vectors with a 1 in $i^{\mathrm{th}}$ position, 0 elsewhere). Additionally, we will often refer to the bases $\mathcal{U}(t)=\{{\bm{u}}_{i}^{(t)}\}_{i=1\dots D}$ formed by the left-singular vectors of the singular value decomposition (SVD) of the $D\times N_{\mathrm{s}}$ matrix after layer $t$: ${\bm{M}}={\bm{U}}\mathbf{\Sigma}{\bm{V}}^{\top}$, with ${\bm{M}}_{:,k}^{(t)}={\bm{x}}_{k}(t)$. Vectors ${\bm{u}}_{i}$ are organized according to their corresponding singular values, $\sigma_{i}$, in descending order. Note that because trajectory clusters evolve over time there are 24 distinct bases.
@@ 3.1 Lines of thought cluster along similar pathways
We endeavor to visualize and characterize the trajectories of pilot tokens in the latent space $\mathcal{S}$. The high dimensionality makes it non-trivial. Which projections should we consider?
There is no reason a priori for the Cartesian axes, ${\bm{e}}_{i}$, to align with any meaningful directions, so we seek relevant alternative bases informed by the data. Naturally, we consider the bases $\mathcal{U}(t)$ formed by the singular vectors of pilots ensemble after each layer $t$.
Using these bases aligned with the data’s intrinsic directions, we observe in Fig. 1a that trajectories tend to cluster together, instead of producing an isotropic and homogeneous filling of $\mathcal{S}$. Indeed, LoTs for different, independent pseudo-sentences follow a common path (forming bundles), augmented by individual variability. Specifically, there exist directions with significant displacement relative to the spread (mean over standard deviation); and positions at different times form distinct clusters, as shown in Fig. A1.
@@ 3.2 Lines of thought follow a low-dimensional manifold
We remark in Fig. 2a that the intrinsic bases $\mathcal{U}(t)$ rotate only slightly across successive timepoints $t$. Besides, Fig. 2b shows that the corresponding singular values decay quickly over several orders of magnitude. Both suggest that LoTs may be described by a low-dimensional curved subspace.
But how many dimensions are relevant? Singular values relate to ensemble variance along their corresponding directions. Since the embedding space is high-dimensional, however, the curse of dimensionality looms, hence the significance of Euclidean distances crumbles. To circumvent this limitation, we consider a more practical metric: how close to the original output distribution on the vocabulary does a reduction in dimensionality get us?
@@
where the ${\bm{u}}_{i}^{(t)}$’s are organized by descending order of their corresponding singular values. By default $K=D$, and the true output distribution ${\bm{p}}^{\mathcal{V}}$ is obtained. Now, we examine what happens when, instead of passing the full basis set, we truncate it, after each layer, to keep only the first $K<D$ principal components. We compare the resulting output distribution, $\mathbf{p}^{\mathcal{V}}_{K}$ to the true distribution $\mathbf{p}^{\mathcal{V}}$ using KL divergence $D_{\mathrm{KL}}(\mathbf{p}^{\mathcal{V}}_{K}\|\mathbf{p}^{\mathcal{V}})$. In Fig. 2c, we see that $D_{\mathrm{KL}}$ decreases very slowly with decreasing $K$, up to about $K_{0}=256$. At that point, $D_{\mathrm{KL}}$ is only about 10% of its uncorrelated baseline value, implying that most of the true distribution is recovered when keeping only about $K_{0}=256$, or $25\%$, of the principal components. In other words, for the purpose of next-token prediction, LoTs are quasi-256-dimensional.
@@
Figure 2: (a) Angle between the first 4 singular vectors at $(t_{1},t_{2})$, $\arccos({\bm{u}}_{i}^{(t_{1})}\cdot{\bm{u}}_{i}^{(t_{2})})$, for $i=\{1,2,3,4\}$ (top-left, top-right, bottom-left, bottom-right, respectively). (b) Singular values for $t=1,\dots,24$ (blue to red). Clusters stretch more and more after each layer. The leading singular values, $\sigma_{1}(t)$, have been omitted for clarity. (c) Average (over all trajectories) KL divergence between reduced dimensionality trajectories output and true output distributions, as the dimensionality $K$ is increased. The red dashes line shows the average KL divergence for output distributions from unrelated inputs
(baseline for dissimilar distributions).
@@ 3.3 Linear approximation of trajectories
Examination of the singular vectors and values at each time step indicates that LoT bundles rotate and stretch smoothly after passing through each layer (Fig. 2). This suggests that token trajectories could be approximated by the linear transformations described by the ensemble, and extrapolated accordingly, from an initial time $t$ to a later time $t+\tau$. Evidently, it is improbable that a transformer layer could be replaced by a mere linear transformation. We rather hypothesize that, in addition to this deterministic average path, a token’s location after layer $t+\tau$ will depart from its linear approximation from $t$ by an unknown component ${\bm{w}}(t,\tau)$.777 We emphasize that prompt trajectories are completely deterministic; the stochastic component introduced in the model accounts for the fact that we perform a linear extrapolation based only on a token’s position at a certain time, which unsurprisingly deviates from the true position obtained from processing the full prompt with transformer layers. We propose the following model:
@@
where ${\bm{x}}(t)$ is the pilot token’s position in the Cartesian basis, and ${\bm{R}},\mathbf{\Lambda}$ are rotation (orthonormal) and stretch (diagonal) matrices, respectively. Eq. 1 formalizes the idea that, to approximate ${\bm{x}}(t+\tau)$, given ${\bm{x}}(t)$, we first project ${\bm{x}}$ in the ensemble intrinsic basis at $t$ (${\bm{R}}^{\top}{\bm{x}}$), then stretch the coordinates by the amount given by $\mathbf{\Lambda}$, and finally rotate according to how much the singular directions have rotated between $t$ and $t+\tau$, ${\bm{R}}(t+\tau)$ (see also Fig. A2 in Appendix B). Consequently, we can express these matrices as a function of the set of singular vectors (${\bm{U}}$) and values ($\mathbf{\Sigma}$):
@@
Fig. 3 shows the close agreement, at the ensemble level, between the true and extrapolated positions.
This is confirmed by the observation that the two sets are not separable with a trained linear classifier, including at large $\tau$ (see Appendix). Eq. 1 is merely a linear approximation as it is similar to assuming that LoT clusters deform like an elastic solid, where each point maintains the same vicinity, as illustrated in Fig. A2. The actual coordinates ought to include an additional random component ${\mathbf{w}}(t,\tau)$, which a priori depends on both $t$ and $\tau$.
@@
between true positions ${\bm{x}}$ and linear approximations $\tilde{{\bm{x}}}(t,\tau)={\bm{R}}(t+\tau)\mathbf{\Lambda}(t,\tau){\bm{R}}(t)^{\top}{\bm{x}}(t)$. We investigate the distributions and correlations of $\delta{\bm{x}}(t,\tau)$ across layer combinations $(t,t+\tau)$.
From the data, Fig. 4 shows that, for all $(t,t+\tau)\in\{1,\dots,23\}\times\{t+1,\dots,24\}$, the ensemble of $\delta{\bm{x}}(t,\tau)$ has the following characteristics: 1) it is Gaussian, 2) with zero mean, 3) and variance scaling as $\exp(t+\tau)$. In addition, Fig. A3 shows that the distribution is isotropic, with no evidence of spatial cross-correlations . Hence, we propose:
@@ 3.4 Langevin dynamics for continuous time trajectories
In return, this allows us to interpolate trajectories between transformer layers.999 These interpolated positions do not hold any interpretive value, but may be insightful for mathematical purposes. Thus, we extend Eq. 1 to a continuous time variable $t$, and write in infinitesimal terms the Langevin equation for the dynamics:
@@ 4.2.1 Gibberish
We generate non-language (‘gibberish’) pseudo-sentences by assembling $N$-token sequences of random tokens in the vocabulary, and pass them as input to GPT-2. The resulting trajectories also cluster around a path similar to that of language. However, the two ensembles, language and gibberish, are linearly separable at all layers (see Fig. A5 in Section B.6), indicating that they travel on two distinct, yet adjacent, manifolds.
@@ 4.2.2 Untrained & ablated models
First, we collect trajectories of the Walden ensemble passing through a reinitialized version of GPT-2 (the weights have been reset to a random seed). We observe that while LoTs get transported away from their starting point, the trajectories follow straight, quasi-parallel paths, maintaining their vicinity (see Fig. A5). Furthermore, the model of Eqs. 1 and 2 does not hold; Fig. A6 shows that the variance of $\delta{\bm{x}}$ does not follow the $\exp(t+\tau)$ scaling, and the distributions are far from Gaussian.
Next, we consider an ablated model, where only layers $13$ to $24$ have been reinitialized. When reaching the untrained layers, the trajectories stop and merely diffuse about their $t=12$ location (Fig. A5).
@@ 4.3 Results with other models
We first investigate the Llama 2 7B model (Touvron et al., 2023).111111 Decoder-only, 32 layers, 4096 dimensions; released July 2023 by Meta AI. Remarkably, the pattern of GPT-2 repeats. Token positions at $t+\tau$ can be extrapolated from $t$ by rotation and stretch using the singular vectors and values of the ensemble. The residuals are distributed as those of GPT-2, with ${\textnormal{w}}_{i}(t,\tau)\sim\mathcal{N}(0,\alpha e^{\lambda(t+\tau)})$, see Fig. A7. The values for the parameters $\alpha$ and $\lambda$, however, differ from those of GPT-2 (here, $\alpha\simeq-5.4,\lambda\simeq 0.27$).
@@
Trajectories across the Mistral 7B (v0.1) model (Jiang et al., 2023)121212 Decoder-only, 32 layers, 4096 dimensions; released September 2023 by Mistral AI. also follow the same pattern (Fig. A8). We note, however, that Eq. 4 only holds up until layer 31. It seems as though the last layer is misaligned with the rest of the trajectories, as linear extrapolation produces an error that is much larger than expected.
@@
The last layer anomaly is also apparent for Llama 3.2 1B131313 Decoder-only, 16 layers, 2048 dimensions; released September 2024 by Meta AI. , both in the mean and variance of $\delta{\bm{x}}(t,16)$ (see Fig. A9). However, the rest of the trajectories follows Eq. 1. The same pattern is observed for Llama 3.2 3B141414 Decoder-only, 28 layers, 3072 dimensions; released September 2024 by Meta AI. in Fig. A10.